A 5.0-liter gas tank holds 1.4 moles of helium (He) and 0.70 moles of oxygen (O2), at a temperature of 260 K.

The atomic masses of helium and oxygen are 4.0 g/mol and 16.0 g/mol, respectively.

The total random translational kinetic energy of the gas in the tank is closest to


1
Expert's answer
2021-08-06T15:26:24-0400

K for one molecule:

"K = \frac{3}{2}kT \\

= \frac{3}{2}(1.38 \times 10^{-23}\;J/K) (260\;K) \\

= 5.38 \times 10^{-21} \;J"

All molecules:

"(1.4 \times 6.023 \times 10^{23} + 0.7 \times 6.023 \times 10^{23}) \times 2.38 \times 10^{-21} \;J \\

= 6.8 \;kJ"